According to a report from the center off studying health system change 20 percent of Americans will delay or go without medical are because of concerns about cost. Suppose 8 individuals are randomly selected. What is the probability that at least 7 will delay or go without medical care. Round answer to 4 decimal places and show work.
1
Expert's answer
2012-10-04T09:34:52-0400
Using Poisson's distribution we calculate P(7) and P(8) using it's formula:

P(7)=(0.2)^7/7!*e^(-0.2)=0.0000
P(8)=(0.2)^8/8!*e^(-0.2)=0.0000
P(x>=7)=P(x=7)+P(x=8)=0.0000